OpenStudy (anonymous):

(y+3)^3+8 ? how to factor completely ?

OpenStudy (usukidoll):

sum of cubes formula \[(a+b)^3 =(a+b)(a^2-ab+b^2) \] rewrite \[(y+3)^3=8\] as \[(y+3)^3=2^3\] so let a = y+3 and b = 2

OpenStudy (anonymous):

hmm... then ?:(

OpenStudy (usukidoll):

well the b portion is easy because if b = 2 then b^2 is ?

OpenStudy (usukidoll):

OH MAN! I GOT MY LATEX WRONG NUGH! \[\large (y+3)^3+ 8 \rightarrow = (y+3)^3+2^3 \]

OpenStudy (usukidoll):

but the sum of cubes formula is correct thank goodness.

OpenStudy (usukidoll):

so if b = 2 what's b^2 = ?

OpenStudy (anonymous):

4?

OpenStudy (usukidoll):

\[\large (y+3)^3+ 8 \rightarrow (y+3)^3+2^3\] idk what's with me in the middle of the night

OpenStudy (usukidoll):

but yes b^2 = 4 because b = 2 2^2 = 4 woo.

OpenStudy (usukidoll):

\[((y+3)+(2))^3 =((y+3)+2)((y+3)^2-2(y+3)+4)\]

OpenStudy (anonymous):

(y+3+2) = (y+5) (_______)

OpenStudy (usukidoll):

so we need to expand (y+3)^2 and distribute the -2 on y+3

OpenStudy (anonymous):

(y+5)(y^2+9-(y+3+2)+4) ?

OpenStudy (usukidoll):

wait.. try expand this in parts expand \[(y+3)^2 \] and then distribute -2(y+3)

imqwerty (imqwerty):

i didn't even say something :(

OpenStudy (anonymous):

y^2+3y+9 then?

imqwerty (imqwerty):

i ws here to help k..sry bye i won't interrupt u frm nw .

OpenStudy (usukidoll):

then distribute the -2 for -2(y+3)

OpenStudy (anonymous):

-2y-6

OpenStudy (usukidoll):

yes so now we can place them into our equation \[((y+3)+(2))^3 =((y+3)+2)((y+3)^2-2(y+3)+4) \] \[((y+3)+(2))^3 =(y+5)(y^2+3y+9-2y-6+4) \]

OpenStudy (usukidoll):

then we can rearrange and combine like terms...

OpenStudy (usukidoll):

\[(y+5)(y^2+3y-2y-6+4+9)\]

OpenStudy (usukidoll):

so combine like terms in the super long equation what's 3y-2y what's -6+4+9

OpenStudy (anonymous):

(y+5)(y^2+y+7) ?

OpenStudy (usukidoll):

cool so now we need to ... wow distribute \[(y+5)(y^2+y+7)\] so distribute the y throughout \[y^2+y+7\] like \[y(y^2+y+7)\]

OpenStudy (usukidoll):

\[(y^2(y)+y(y)+7(y))\]

OpenStudy (usukidoll):

do you know the exponent laws ?

OpenStudy (anonymous):

yup

OpenStudy (usukidoll):

alrighty then :)

OpenStudy (anonymous):

y^3+y^2+7y

OpenStudy (usukidoll):

excellent! now we do the same thing for the 5 \[(y+5)(y^2+y+7) \] so distribute 5 throughout \[y^2+y+7\] \[5(y^2+y+7)\]

OpenStudy (usukidoll):

so it's like 5 x y^2 5 x y 5 x 7

OpenStudy (anonymous):

5y^2 +5y+35

OpenStudy (usukidoll):

yeah! so now we combine like terms \[y^3+y^2+7y+5y^2+5y+35\]

OpenStudy (usukidoll):

-_- something is wrong ... our middle terms are off.

OpenStudy (anonymous):

y^3+5y^2+12y+35

OpenStudy (usukidoll):

yeah but I'm double checking through a calculator and it claims that it's 9y^2+27y unless ... hey @Haseeb96 could you come over here and check this?

OpenStudy (usukidoll):

oh nevermind.. wait let me think... our first term and last term is correct..

OpenStudy (anonymous):

i hate factoring completely :(

OpenStudy (anonymous):

bye :) i have lot to do :/

OpenStudy (usukidoll):

ok. I'll try and figure out what's up before I go sleep. Maybe we needed pascal's triangle/binomial theorem.

OpenStudy (anonymous):

thanks for your time dude :)

OpenStudy (usukidoll):

you're welcome. I'm gonna post the pascal's triangle version.. I see it clearer on how they got the middle term... another method is to exapand (y+3)^3 manually and add 8 but who wants to do that ? that's time consuming.

OpenStudy (usukidoll):